OpenStudy (anonymous):

Determine whether or not the function f:ZxZ->Z is onto, if f((m,n))=mn. Determine whether or not the function f:ZxZ->Z is onto, if f((m,n))=mn. @Mathematics

OpenStudy (anonymous):

And what does ZxZ->Z mean?

OpenStudy (jamesj):

a function from the 2-dim lattice of integers to the integers.

OpenStudy (jamesj):

Z = the integers = {0, +1, -1, +2, -2, +3, -3, ....}

OpenStudy (jamesj):

and ZxZ is the cartesian product of Z with itself, which you can think of as a lattice of points on the usual 2-dimensional plane.

OpenStudy (jamesj):

Now, if you understand that and you know the definition of onto, it should be easy for you to determine whether this function f is onto or not.

OpenStudy (anonymous):

so you're saying it's a function that can point to more than one thing?

OpenStudy (jamesj):

What is the formal definition of this statement a function f: X -> Y is ONTO if ....what?

OpenStudy (anonymous):

I actually don't know haven't learnt onto yet

OpenStudy (anonymous):

I'll look it up give me a sec

OpenStudy (jamesj):

It means that for every \( y \in Y \) there exists an \( x \in X \) such that \(f(x) = y\). On the real numbers, the function f(x) = x is onto (and also 1:1). The function f(x) = x^3 is also onto. The function f(x) = x^3 - x is also onto, but not 1:1. The function f(x) = sin(x) is not onto because for y = 2 for instance, there is no x such that sin(x) = 2.

OpenStudy (jamesj):

The function \( f : R^3 \rightarrow R^2 \) given by \[ f(x,y,z) = (x,y) \] is also onto but not 1:1

OpenStudy (anonymous):

Oh I see I still don't get what zxz->z means so all integers including 0?

OpenStudy (jamesj):

\( \mathbb{Z} \) is the standard symbol in mathematics for the integers, including negatives and zero.

OpenStudy (anonymous):

okay but why does it do zxz it's saying it's 2 dimensional?

OpenStudy (anonymous):

Is it onto because for every number if we make m=k and n=1 we'll get k?

OpenStudy (jamesj):

\( A \times B \) is the Cartesian product of sets, \[ A \times B = \{ (a,b) \ | \ a \in A, b \in B \} \] This is why we write \( \mathbb{R}^2 \) for the two-dimensional Cartesian plane: \[ \mathbb{R}^2 = \mathbb{R} \times \mathbb{R} \] Similarly, \[ \mathbb{R}^3 = \mathbb{R} \times \mathbb{R} \times \mathbb{R} \]

OpenStudy (jamesj):

Thus \[ \mathbb{Z} \times \mathbb{Z} = \{ (p,q) \ | \ p, q \in \mathbb{Z} \} \]

OpenStudy (jamesj):

Now, "Is it onto because for every number if we make m=k and n=1 we'll get k?" yes. Just formalize that argument so it makes exactly the definition. So the argument starts like this. "Given an arbitrary integer \( p \in Z \) let (m,n) = ( ...), then f(m,n) = p. Hence, by definition, f is onto Z"

OpenStudy (anonymous):

Thank you!! That makes a lot more sense!